# MVP-ORAM: a Wait-free Concurrent ORAM for Confidential BFT Storage

### v1

@@ Abstract
It is well known that encryption alone is not enough to protect data privacy. Access patterns, revealed when operations are performed, can also be leveraged in inference attacks. Oblivious RAM (ORAM) hides access patterns by making client requests oblivious. However, existing protocols are still limited in supporting concurrent clients and Byzantine fault tolerance (BFT). We present MVP-ORAM, the first wait-free ORAM protocol that supports concurrent fail-prone clients. In contrast to previous works, MVP-ORAM avoids using trusted proxies, which require additional security assumptions, and concurrency control mechanisms based on inter-client communication or distributed locks, which limit overall throughput and the capability of tolerating faulty clients. Instead, MVP-ORAM enables clients to perform concurrent requests and merge conflicting updates as they happen, satisfying wait-freedom, i.e., clients make progress independently of the performance or failures of other clients. Since wait and collision freedom are fundamentally contradictory goals that cannot be achieved simultaneously in an asynchronous concurrent ORAM service, we define a weaker notion of obliviousness that depends on the application workload and number of concurrent clients, and prove MVP-ORAM is secure in practical scenarios where clients perform skewed block accesses. By being wait-free, MVP-ORAM can be seamlessly integrated into existing confidential BFT data stores, creating the first BFT ORAM construction. We implement MVP-ORAM on top of a confidential BFT data store and show our prototype can process hundreds of 4kB accesses per second in modern clouds.
publicationid: pubid: This is the extended version of the paper appearing at the 33rd Network and Distributed System Security Symposium (NDSS 2026).
@@ I Introduction
Context and motivation. Byzantine Fault-Tolerant State Machine Replication (BFT SMR) is a classical technique to implement fault- and intrusion-tolerant replicated services with strong consistency [schneider1990implementing, castro1999practical]. The technique attracted significant attention in the last decade due to the emergence of decentralized systems and blockchains [nakamoto_2009, wood_2014], which can be seen as replicated state machines. BFT SMR systems offer data integrity and availability guarantees, even if a fraction $t$ of the $n$ replicas is compromised. Some works have additionally studied how to make BFT SMR systems offer confidentiality  [depspace, padilha_2011, basu_2019, calypso, vassantlal_2022], guaranteeing data secrecy even in the presence of Byzantine faults. This is typically achieved by combining symmetric encryption with secret sharing [shamir1979share, blakely1979safeguarding], a technique where a secret (e.g., an encryption key) is split into $n$ shares (one for each server) and any subset of $t+1$ of them is required for recovering it.
@@
State of the art. Oblivious RAM (ORAM) [goldreich1987towards, ostrovsky1990efficient, goldreich1996software] is a cryptographic technique whose objective is to conceal access patterns and make them oblivious, i.e., indistinguishable from each other. However, it was initially designed for a single CPU accessing its RAM (or a client accessing its server). Recent works have studied its suitability to support multiple CPUs/clients in different yet related lines of work known as parallel ORAM [boyle2015, chen2016oblivious, hubert2017circuit, chan2018perfectly, chan2020perfectly, asharov2022optimal] and multi-client ORAM [goodrich2012privacy, sahin2016taostore, blass2017multi, maffei2017maliciously, concuroram]. However, these works typically require inter-client communication for synchronization [asharov2022optimal], distributed locks [concuroram], or trusted proxies/hardware [sahin2016taostore, sasy2017zerotrace], all of which severely limit concurrency or require additional security assumptions. Concurrently, researchers have also studied how multiple servers can be leveraged in ORAM [stefanov2013multi, s3oram, larsen2020lower, hoang2020multi, macao] to reduce client-server bandwidth. To the best of our knowledge, QuORAM [maiyya2022quoram] is the only multi-server protocol that increases ORAM availability but only addresses server crashes. Hence, to this date, no ORAM protocol can conceal access patterns in Byzantine-resilient systems, and existing protocols can not be extended to meet the requirements of BFT SMR without sacrificing client fault tolerance.
Problem statement. We address the problem of designing practical BFT-replicated storage services that can hide data access patterns, going beyond existing works on confidential BFT systems (e.g., [depspace, padilha_2011, basu_2019, calypso]). We aim to design a replicated ORAM service in which replicas can be subject to Byzantine failures while concurrently accessed by fail-prone clients. In more detail, we are interested in a setting where concurrent clients access a replicated storage service in which (1) servers are subject to Byzantine faults, (2) multiple clients can concurrently access shared data stored in servers without external coordination or waiting for each other, and (3) data access patterns remain oblivious. In this setting, a particularly important property to achieve is wait-freedom [herlihy_1991], which states that every client operation (ORAM access) is guaranteed to finish in a finite number of steps. We stress the practical relevance of this property, as wait-free services are more robust than lock-based services [ZooKeeper], which are the norm in concurrent ORAM literature, as clients can finish their accesses independently of the delays and faults of other clients.
@@
MVP-ORAM improves Path ORAM in two fundamental ways. First and foremost, it supports multiple clients concurrently accessing data while satisfying strong consistency [herlihy_1990]. Contrary to previous works on parallel ORAM, where locks and other inter-client coordination are used to ensure consistency and security, we aim to support client-independent, wait-free [herlihy_1991] ORAM accesses, as required in SMR-based services. To the best of our knowledge, no ORAM satisfies this property.
@@
Figure 1: Multiple versions being merged in MVP-ORAM.
@@
To make our oblivious BFT data store feature-complete, we tackle the problem of managing the shared keys used to encrypt server data. Contrary to previous works on the multi-client setting, which assume that encryption keys are shared between clients in some way, MVP-ORAM integrates a secret-sharing framework to distribute encryption keys through the servers alongside the ORAM state [basu_2019, vassantlal_2022]. Fig. 2 illustrates our construction, where MVP-ORAM runs on top of a secret sharing and BFT SMR.
@@
We implemented MVP-ORAM on top of COBRA [vassantlal_2022], an open-source confidential BFT SMR library, and evaluated its performance. Our results show that our prototype can process more than $350$ (resp. $700$) accesses per second with a latency of less than $140$ ms (resp. $70$ ms) in a system with $10$ servers (resp. a single server) and $50$ concurrent clients. This shows MVP-ORAM can achieve performance numbers in line with other practical (concurrent/fault-tolerant) protocols [concuroram, maiyya2022quoram].
@@
    We use MVP-ORAM to hide access patterns of a confidential BFT SMR-based storage service, providing the first ORAM that tolerates Byzantine-faulty serversVI);
@@ II Background and Related Work
The research gap: BFT ORAM. As evidenced by our previous discussion, no ORAM protocol can be integrated “as is” into a confidential BFT system to hide access patterns. The state-of-the-art in ORAM fault tolerance is QuORAM [maiyya2022quoram], but it only tolerates crashes and requires trusted execution support on servers, i.e., each server needs a trusted proxy. Extending it to tolerate Byzantine faults seems doable, but it would still require trusted proxies. Regarding multi-client support without using proxies, ConcurORAM [concuroram] is the state-of-the-art. However, it has three main limitations. First, it heavily relies on multi-threading and locks at the server, which introduces nondeterminism, significantly complicating replication (e.g., [locks]). Second, wait-freedom is impossible to achieve using locks, seriously compromising client fault tolerance. Third, it requires 18 server requests to complete one ORAM access (query+eviction), which, if replicated, would require Byzantine consensus for totally ordering each request. This large number of client-server iterations is also a limitation of a recent Optimal parallel ORAM by Asharov et al. [asharov2022optimal]. In contrast, by extending Path ORAM (which only requires two round-trips per access) to keep multiple versions of the ORAM state, MVP-ORAM supports concurrent clients without using locks and requiring just three round-trips per ORAM access (see Fig. 3), making it thus more appropriate to be integrated into confidential BFT SMR systems.
@@ III Model and Definitions
Service model. Clients access the replicated storage service, which contains $N$ data blocks, by sending requests and receiving replies to/from the service replicas. Servers globally store a two-part state $S=\langle C,P\rangle$. The common state $C$ comprises ORAM data, and the private state $P$ comprises encryption keys used to encrypt $C$. Each server $r_{i}$ locally maintains a state $S_{i}=\langle C,P_{i}\rangle$. The common state $C$ is encrypted and replicated across all servers, i.e., all servers store the same state, while the private state is distributed using the secret sharing protocol. Hence, $r_{i}$’s private state $P_{i}$ comprises shares of the encryption keys. The functionality of our ORAM service offers a single operation, described in the following way:
@@
This definition is weaker than the one used in traditional parallel ORAM (e.g., [boyle2015, chen2016oblivious, chan2017]), as it does not allow the adversary to inject queries and it depends on additional security parameters that may make $\mu$ non-negligible, namely $c$ and $\mathcal{D}$. In MVP-ORAM, $\mu$ will be negligible if, per timestep, a single client accesses the ORAM or multiple clients access different blocks. However, if multiple clients access the same block concurrently, $\mu$ may not be necessarily negligible, although it can be arbitrarily small (see §VII-A for more details).
@@ IV-B Extending Path ORAM to Multiple Clients
  • $\mathit{pm}\leftarrow\mathsf{getPM}(c_{i})$: if there is no active client, sets client $c_{i}$ as active (i.e., started an access operation) and retrieves the current position map $\mathit{pm}$. Else, it returns $\bot$.
  • $\langle\mathcal{P}_{l},S\rangle\leftarrow\mathsf{getPS}(c_{i},l)$: if $c_{i}$ is still the active client, the server returns path $\mathcal{P}_{l}$ and stash $S$; otherwise the server returns $\bot$.
@@
Although each individual step is atomic at the server, the access is not, as it requires three steps, and different clients can interleave these steps, interrupting accesses one from another. As a result, this first protocol only ensures a client completes its access if no other client accesses the ORAM concurrently, satisfying obstruction freedom [herlihy_2003].
@@ V Multi-Version Path ORAM
To further increase the number of available paths, we keep the last accessed block in the stash, meaning it can be accessed again using any path, and evict the most frequently accessed blocks to higher levels of the tree. Specifically, when performing an access, the accessed block always goes to the stash (if it is not already there) together with the non-dummy blocks from $Z$ slots uniformly selected at random. Then, these $Z$ slots are filled with $Z$ random blocks previously in the stash, i.e., we swap $Z$ blocks between the stash and the accessed path. The constant $Z$ is important to bound the stash size. Additionally, after the swap, we reorder blocks in the path by placing the most frequently accessed blocks higher in the tree. The result is that after each access, the most popular blocks in a skewed access pattern will be accessible through many paths, improving the ORAM obliviouness.
To tackle the second problem, we allow clients to finish their access in isolation and store updates as new versions of the tree on the server. During an access, each client fetches the existing versions currently stored in the server and merges them into a single, updated tree (as illustrated in Fig. 1). Note that in practice, clients only need to merge paths and stashes that they will retrieve in an access, instead of whole trees.
When multiple versions accessed by clients are merged together, such a merge needs to be done (1) without losing block updates, [^2] Linerizability [herlihy_1990], or register atomicity, requires a read executed after a write to always see the last update on the stored data. (2) by keeping more frequently accessed blocks on higher levels of the tree, and (3) by avoiding block duplication in different tree nodes during concurrent evictions. To satisfy these requirements, the server assigns a sequence number to each access during the invocation of its $\mathsf{Server.getPM}$. This sequence number is used to create a logical block timestamp $\mathit{ts}_{b}=\langle v,a,s\rangle$ for each block $b$ touched during an access, where version $v$ is the sequence number of the last write on this block, access $a$ is the sequence number of the last read or write of this block, and sequence $s$ is the sequence number of the last time the block was moved. For instance, if the sequence number of an access to block $b$ is $x$, $\mathit{ts}_{b}$ after the operation will be $\langle x,x,x\rangle$ if $b$ is written or $\mathit{ts}_{b}=\langle\_,x,x\rangle$ if $b$ is read. Further, any other block that changed its slot during this access’ eviction will have its timestamp set to $\langle\_,\_,x\rangle$.
@@ 21 $\mathcal{S}^{*}\leftarrow(\mathcal{S}^{c}\setminus\mathcal{S})\cup\{S\}$
When the server receives the request, it stores a reference to the current ORAM state in $c_{i}$’s context until the client completes its access. Then, it returns the path map history $\mathcal{H}_{\mathit{pathMaps}}$, and the op. sequence number $\mathit{seq}$ (A2, L4-L8).
After receiving $\mathcal{H}_{\mathit{pathMaps}}$, $c_{i}$ consolidates it into position map $\mathit{pm}$ by invoking $\underline{\mathit{consolidatePathMaps}}$ (First Step of Fig. 4). This function applies the updates contained in $\mathcal{H}_{\mathit{pathMaps}}$ to the local position map by retaining, for each block, the location update with the highest timestamp. With the updated position map $\mathit{pm}$ containing the most recent locations of the blocks, $c_{i}$ discovers the current slot $\mathit{sl}$ where $b$ is stored. The first step terminates with $c_{i}$ randomly choosing a path $\mathcal{P}_{l}$ that contains $\mathit{sl}$ (A1, L5).

### v2

@@ Abstract
It is well known that encryption alone is not enough to protect data privacy. Access patterns, revealed when operations are performed, can also be leveraged in inference attacks. Oblivious RAM (ORAM) hides access patterns by making client requests oblivious. However, existing protocols are still limited in supporting concurrent clients and Byzantine fault tolerance (BFT). We present MVP-ORAM, the first wait-free ORAM protocol that supports concurrent fail-prone clients. In contrast to previous works, MVP-ORAM avoids using trusted proxies, which necessitate additional security assumptions, and concurrency control mechanisms based on inter-client communication or distributed locks, which limit overall throughput and the capability to tolerate faulty clients. Instead, MVP-ORAM enables clients to perform concurrent requests and merge conflicting updates as they happen, satisfying wait-freedom, i.e., clients make progress independently of the performance or failures of other clients. Since wait and collision freedom are fundamentally contradictory goals that cannot be achieved simultaneously in an asynchronous concurrent ORAM service, we define a weaker notion of obliviousness that depends on the application workload and number of concurrent clients, and prove MVP-ORAM is secure in practical scenarios where clients perform skewed block accesses. By being wait-free, MVP-ORAM can be seamlessly integrated into existing confidential BFT data stores, creating the first BFT ORAM construction. We implement MVP-ORAM on top of a confidential BFT data store and show our prototype can process hundreds of 4KB accesses per second in modern clouds.
publicationid: pubid: This is the extended version of the paper published at the Proc. of the 33rd Network and Distributed System Security Symposium (NDSS 2026) [mvp_oram_ndss].
@@ I Introduction
Context and motivation. Byzantine Fault-Tolerant State Machine Replication (BFT SMR) is a classical technique to implement fault- and intrusion-tolerant replicated services with strong consistency [schneider1990implementing, castro1999practical]. The technique attracted significant attention in the last decade due to the emergence of decentralized systems and blockchains [nakamoto_2009, wood_2014], which can be seen as replicated state machines. BFT SMR systems offer data integrity and availability guarantees, even if up to $t$ of the $n$ replicas are compromised. Some works have additionally studied how to make BFT SMR systems offer confidentiality  [depspace, padilha_2011, basu_2019, calypso, vassantlal_2022], guaranteeing data secrecy even in the presence of Byzantine faults. This is typically achieved by combining symmetric encryption with secret sharing [shamir1979share, blakely1979safeguarding], a technique where a secret (e.g., an encryption key) is split into $n$ shares (one for each server) and any subset of $t+1$ of them is required for recovering it.
@@
State of the art. Oblivious RAM (ORAM) [goldreich1987towards, ostrovsky1990efficient, goldreich1996software] is a cryptographic technique whose objective is to conceal access patterns and make them oblivious, i.e., indistinguishable from each other. However, it was initially designed for a single CPU accessing its RAM (or a client accessing its server). Recent works have studied its suitability for supporting multiple CPUs/clients in different yet related lines of work, known as parallel ORAM [boyle2015, chen2016oblivious, hubert2017circuit, chan2018perfectly, chan2020perfectly, asharov2022optimal] and multi-client ORAM [goodrich2012privacy, sahin2016taostore, blass2017multi, maffei2017maliciously, concuroram]. However, these works typically require inter-client communication for synchronization [asharov2022optimal], distributed locks [concuroram], or trusted proxies/hardware [sahin2016taostore], all of which severely limit concurrency or require additional security assumptions. Concurrently, researchers have also studied how multiple servers can be leveraged in ORAM [stefanov2013multi, s3oram, larsen2020lower, hoang2020multi, macao] to reduce client-server bandwidth. To the best of our knowledge, QuORAM [maiyya2022quoram] is the only multi-server protocol that increases ORAM availability but only addresses server crashes. Hence, to date, no ORAM protocol can conceal access patterns in Byzantine-resilient systems, and existing protocols cannot be extended to meet the requirements of BFT SMR without sacrificing client fault tolerance.
Problem statement. We address the problem of designing practical BFT-replicated storage services that can hide data access patterns, going beyond existing works on confidential BFT systems (e.g., [depspace, padilha_2011, basu_2019, calypso, vassantlal_2022]). We aim to design a replicated ORAM service in which replicas can be subject to Byzantine failures while concurrently accessed by fail-prone clients. In more detail, we are interested in a setting where concurrent clients access a replicated storage service in which (1) servers are subject to Byzantine faults, (2) multiple clients can concurrently access shared data stored in servers without external coordination or waiting for each other, and (3) data access patterns remain oblivious. In this setting, a particularly important property to achieve is wait-freedom [herlihy_1991], which states that every client operation (ORAM access) is guaranteed to finish in a finite number of steps. We stress the practical relevance of this property, as wait-free services are more robust than lock-based services [ZooKeeper] (the norm in the concurrent ORAM literature), as clients can finish their accesses independently of the delays and faults of other clients.
@@
MVP-ORAM improves Path ORAM in two fundamental ways. First and foremost, it supports multiple clients concurrently accessing data while satisfying strong consistency [herlihy_1990]. Contrary to previous works on parallel ORAM, which use locks and other inter-client coordination mechanisms to ensure consistency and security, we aim to support client-independent, wait-free ORAM accesses, as required in SMR-based services [castro1999practical]. To the best of our knowledge, no ORAM satisfies this property.
@@
Figure 1: Multiple versions being created by concurrent accesses and later merged in MVP-ORAM.
@@
To make our oblivious BFT data store feature-complete, we tackle the problem of managing the shared keys used to encrypt server data. Contrary to previous works on the multi-client setting, which assume that encryption keys are shared between clients in some way, MVP-ORAM integrates a secret-sharing framework to distribute encryption keys through the servers alongside the ORAM state [basu_2019, vassantlal_2022]. Fig. 2 illustrates our construction, where MVP-ORAM runs on top of secret sharing and BFT SMR.
@@
We implemented MVP-ORAM on top of COBRA [vassantlal_2022], an open-source confidential BFT SMR library, and evaluated its performance. Our results show that our prototype can process more than $350$ (resp. $700$) accesses per second with a latency of less than $140$ ms (resp. $70$ ms) in a system with $10$ servers (resp. a single server) and $50$ concurrent clients. This shows MVP-ORAM can achieve performance numbers in line with other practical (concurrent/fault-tolerant) protocols [concuroram, maiyya2022quoram]. Our implementation is open source [mvp_oram], and our results are fully reproducible, as described in Appendix F.
@@
    We use MVP-ORAM to hide access patterns of a confidential BFT SMR-based storage service, providing the first ORAM that tolerates Byzantine-faulty servers (§VI);
@@ II Background and Related Work
The research gap: BFT ORAM. As evidenced by our previous discussion, no ORAM protocol can be integrated “as is” into a confidential BFT system to hide access patterns. The state-of-the-art in ORAM fault tolerance is QuORAM [maiyya2022quoram], but it only tolerates crashes and requires trusted execution support on servers, i.e., each server needs a trusted proxy. Extending it to tolerate Byzantine faults seems doable, but it would still require trusted proxies. Regarding multi-client support without using proxies, ConcurORAM [concuroram] is the state-of-the-art. However, it has three main limitations. First, it heavily relies on multi-threading and locks at the server, which introduces nondeterminism, significantly complicating replication (e.g., [locks]). Second, wait-freedom is impossible to achieve using locks, seriously compromising client fault tolerance. Third, it requires 18 client-server interactions to complete one ORAM access (query and eviction), which, if replicated, would require Byzantine consensus for totally ordering each request. This large number of client-server iterations is also a limitation of a recent optimal parallel ORAM by Asharov et al. [asharov2022optimal]. In contrast, by extending Path ORAM (which only requires two round-trips per access) to keep multiple versions of the ORAM state, MVP-ORAM supports concurrent clients without using locks and requiring just three round-trips per ORAM access (see Fig. 3), making it thus more appropriate to be integrated into confidential BFT SMR systems.
@@ III Model and Definitions
Service model. Clients access the replicated storage service, which contains $N$ data blocks, by sending requests and receiving replies to/from the service replicas. Servers globally store a two-part state $\Omega=\langle C,P\rangle$. The common state $C$ comprises ORAM data, and the private state $P$ comprises encryption keys used to encrypt $C$. Each server $r_{i}$ locally maintains a state $\Omega_{i}=\langle C,P_{i}\rangle$. The common state $C$ is encrypted and replicated across all servers, i.e., all servers store the same state, while the private state is distributed using the secret sharing protocol. Hence, $r_{i}$’s private state $P_{i}$ comprises shares of the encryption keys. The functionality of our ORAM service offers a single operation, described in the following way:
@@
This definition is weaker than the one used in traditional parallel ORAM (e.g., [boyle2015, chen2016oblivious, chan2017]), as it does not allow the adversary to inject queries and it depends on additional security parameters that may make $\mu$ non-negligible, namely $c$ and $\mathcal{D}$. In MVP-ORAM, $\mu$ will be negligible if, per timestep, a single client accesses the ORAM or multiple clients access different blocks. However, if multiple clients access the same block concurrently, $\mu$ may not be necessarily negligible, although it can be arbitrarily small (see §VII-A).
@@ IV-B Extending Path ORAM to Multiple Clients
  • $\mathit{pm}\leftarrow\mathsf{getPM}(c_{i})$: if there is no active client, sets client $c_{i}$ as active (i.e., started an access operation) and retrieves the current position map $\mathit{pm}$; else, it returns $\bot$.
  • $\langle\mathcal{P}_{l},S\rangle\leftarrow\mathsf{getPS}(c_{i},l)$: if $c_{i}$ is still the active client, the server returns path $\mathcal{P}_{l}$ and stash $S$; otherwise, the server returns $\bot$.
@@
Although each individual step is atomic at the server, the access is not, as it requires three steps, and different clients can interleave these steps, interrupting accesses one from another. As a result, this first protocol only ensures a client completes its access if no other client accesses the ORAM concurrently, satisfying obstruction-freedom [herlihy_2003]. Another consequence of this design is that a failure of a client during an access, which will never end, might block other clients forever.
@@ V Multi-Version Path ORAM
To further increase the number of available paths, we keep the last accessed block in the stash, meaning it can be accessed again using any path, and evict the most frequently accessed blocks to higher levels of the tree. Specifically, when performing an access, the accessed block always goes to the stash (if it is not already there) along with the non-dummy blocks from $Z$ slots uniformly selected at random. Then, these $Z$ slots are filled with $Z$ random blocks previously in the stash, i.e., we swap $Z$ blocks between the stash and the accessed path. The constant $Z$ is important to bound the stash size. Additionally, after the swap, we reorder blocks in the path by placing the most frequently accessed blocks higher in the tree. The result is that after each access, the most popular blocks in a skewed access pattern will be accessible through many paths, improving the ORAM obliviousness.
To address the second problem, we enable clients to complete their access in isolation and store updates as new versions of the tree on the server. During an access, each client fetches the existing versions currently stored in the server and merges them into a single, updated tree (as illustrated in Fig. 1). Note that in practice, clients only need to merge paths and stashes that they will retrieve in an access, rather than entire trees.
When multiple versions accessed by clients are merged together, such a merge needs to be done (1) without losing block updates, [^2] Linerizability [herlihy_1990], or register atomicity, requires a read executed after a write to always return the last update on the stored data. (2) by keeping more frequently accessed blocks on higher levels of the tree, and (3) by avoiding block duplication in different tree nodes during concurrent evictions. To satisfy these requirements, the server assigns a sequence number to each access during the invocation of its $\mathsf{Server.getPM}$. This sequence number is used to create a logical block timestamp $\mathit{ts}_{b}=\langle v,a,s\rangle$ for each block $b$ touched during an access, where version $v$ is the sequence number of the last write on this block, access $a$ is the sequence number of the last read or write of this block, and sequence $s$ is the sequence number of the last time the block was moved. For instance, if the sequence number of an access to block $b$ is $x$, $\mathit{ts}_{b}$ after the operation will be $\langle x,x,x\rangle$ if $b$ is written or $\mathit{ts}_{b}=\langle\_,x,x\rangle$ if $b$ is read. Further, any other block that changed its slot during this access’ eviction will have its timestamp set to $\langle\_,\_,x\rangle$.
@@ 21 $\mathcal{S}^{*}\leftarrow(\mathcal{S}^{c}\setminus\mathcal{S})\cup\{S\}$
When the server receives the request, it stores a reference to the current ORAM state in $c_{i}$’s context until the client completes its access. Then, it returns the path map history $\mathcal{H}_{\mathit{pathMaps}}$, and the access sequence number $\mathit{seq}$ (A2, L4-L8).
After receiving $\mathcal{H}_{\mathit{pathMaps}}$, $c_{i}$ consolidates it into position map $\mathit{pm}$ by invoking $\underline{\mathit{consolidatePathMaps}}$ (first step of Fig. 4). This function applies the updates contained in $\mathcal{H}_{\mathit{pathMaps}}$ to the local position map, retaining for each block the location update with the highest timestamp. With the updated position map $\mathit{pm}$ containing the most recent locations of the blocks, $c_{i}$ discovers the current slot $\mathit{sl}$ where $b$ is stored. The first step terminates with $c_{i}$ randomly choosing a path $\mathcal{P}_{l}$ that contains $\mathit{sl}$ (A1, L5).